# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v2

@@ 2.1.3 Analysis of the approximation and bounds on average sample sizes
The ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ obtained from simulation is represented in Figure 8. As can be seen, it is in general close to $\lambda$, and very close for small or moderate values of $A$. The relatively larger deviations in the rightmost region of the graphs can again be attributed to the fact that for large $A$, since $r_{1}$ is limited from below by $3$ in (81), the value of $r_{1}$ does not result in $c(A,r_{1},\xi)$ close to 0. In any case, the deviation between the actual ratio and $\lambda$ is less than $11\%$ for all combinations of parameters shown in the figure, and much lower for $A$ small.
@@ 2.1.4 Estimation efficiency
This matrix is readily computed from (83) as $F_{i,i}=n_{i}/(p_{i}(1-p_{i}))$, $F_{i,j}=0$ for $i\neq j$ (which reflects the fact that observations from one population give no information about the other), and thus (84) becomes
@@
Figure 9 shows the efficiency $\eta_{\mathrm{el}}$ obtained from Monte Carlo simulation. Its value is computed using (88) with $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ replaced by the corresponding sample means and $\operatorname{Var}[\hat{\theta}$] replaced by the sample MSE. The bound (89) is also plotted. As seen in the figure, the efficiency is high for the values of the target $A$ commonly used in practice, and in particular it is close to $1$ for small $A$. For example, $A=0.04$, corresponding to a relative RMSE of $20\%$, gives values of $\eta_{\mathrm{el}}$ around $80\%$. The efficiency as a function of $A$ exhibits the same discontinuities and steep changes that have been identified for the relative MSE and for $\operatorname{E}[M_{i}+N_{i}]$. The simulation results deviate more from the theoretical bound for large $A$ (rightmost part of the graphs) or for large $\phi$ (Figure 9(b)). This is explained by the fact that in these conditions the relative MSE is considerably smaller than the target, and the average sample sizes are also considerably smaller than their bounds, as discussed previously. On the other hand, the bound is quite tight for small or moderate values of $A$ and $\phi$, which is precisely when it is most important to characterize $\eta_{\mathrm{el}}$ accurately, as sample sizes are large in that case.
It is easily seen that $\lim_{A\rightarrow 0}Ar_{1}^{2}=1$ for $A$ and $r_{1}$ related by (82), and thus also for $r_{1}$ obtained from (81). Combining this result with (89), it follows that $\eta_{\mathrm{el}}$ tends to $1$ when $A$ and $\phi$ tend to 0. This is in consonance with the values shown in Figure 9.
@@ 2.2.1 Estimation procedure
Estimation with group sampling thus proceeds as in Section 2.1, but with an added “outer layer” that translates group sampling into element sampling as described. With the above definition of $M_{i}$ and $N_{i}$, it follows that the number of required groups is
@@ 2.2.2 Average number of groups
and noting that $\max\{x,y\}=(x+y)/2+|x-y|/$2, $\operatorname{E}[G]$ can be expressed as
@@
Using approximations (37) and (38) for $s_{1},s_{2}$ and including the rounding term $\xi$ as in Section 2.1.2 (this has very little effect on the approximations obtained here, but it is done for consistency), (95) becomes
@@
Figure 10 represents the normalized average number of groups obtained from simulation, as well as its theoretical approximation (109). Overall, the graphs follow similar patterns to those observed for $\operatorname{E}[M_{i}+N_{i}]\phi$ with element sampling. The approximation is accurate for small or moderate $A$, and less so for large $A$. Discontinuities due to the choice of $r_{1}$ are minimally visible, and steep changes due to rounding $s_{1}$ and $s_{2}$ can be observed in the simulation results for large $A$. Comparing $\operatorname{E}[G]$ for different values of $\theta$ only makes sense for $l_{1}$ and $l_{2}$ fixed; that is, between Figures 10(a) and 10(b) or between Figures 10(c) and 10(d). For $l_{1}$ and $l_{2}$ given, it can be seen from (109) that the minimum of $\operatorname{E}[G]$ with respect to $\theta$ does not necessarily occur when $\lambda\theta=1$, as was the case for element sampling.
@@ 2.2.3 Estimation efficiency
The efficiency of the estimator is defined, as in Section 2.1.4, by comparing the estimation variance with the lowest variance that could be achieved by a fixed-size estimator with the same average numbers of samples, as given by the Cramér–Rao bound. With group sampling, the average number of samples required from population $i$ is $\operatorname{E}[G]l_{i}$. Equating $n_{i}$ to $\operatorname{E}[G]l_{i}$ in (86), the efficiency $\eta_{\mathrm{gr}}$ for an unbiased estimator of a generic parameter $\zeta$ with group sampling is obtained as
@@
Simulation results for the efficiency are shown in Figure 11. The values are obtained using (111) with $\operatorname{E}[G]$ and $\operatorname{Var}[\hat{\theta}$] replaced by the sample mean and sample MSE. The approximation (112) is also displayed. The figure contains only two specific cases for brevity. The results are in general little sensitive to $\lambda$ and $\theta$; but, similarly to what was observed with element sampling, the theoretical approximation becomes more conservative for large $A$ or $\phi$. Comparing with Figure 9, group sampling is seen to be less efficient than element sampling, in accordance with the number of required groups being the maximum over the two populations, as seen in (91). This causes an efficiency loss of approximately $0.15$ for values of $A$ in the range 0.010.010.10.1, and a less substantial loss for $A$ small.
@@ 3 Estimation of log relative risk
This section describes the LRR estimator and analyzes its properties. The presentation, as will also be the case for subsequent estimators, follows the same logical course as in Section 2, but can be shorter, thanks to the similarities with RR. Section 3.1 addresses the estimation procedure with element sampling, and Section 3.2 considers group sampling.
@@ 3.1 Element sampling
$\operatorname{Var}[\hat{\Theta}^{2}]=\operatorname{E}[\operatorname{E}[\hat{\Theta}^{2}\mid s_{1},s_{2}]]-\Theta^{2}<\operatorname{E}\left[e(s_{1},s_{2})+\Theta^{2}\right]-\Theta^{2}<A.$ (117)
@@
As in the RR case, $s_{1}$ and $s_{2}$ are obtained from the equation system formed by (20) and (33), except with $\mu_{1},\mu_{2},\mu_{12}$ given by (116). The solution is (64)–(66). The values of $s_{1}$ and $s_{2}$ should then be rounded while maintaining (31). Using the first-order approximations (37)–(42), $\gamma$ and $\delta_{1}$ are obtained as as in (61) and (62); and (22) is satisfied. The average sample sizes are approximately given by (70) and (71).
The curvature function is defined as for RR, and the condition $c(A,r_{1},\xi)=0$ is expressed by (82). The fact that $\mu_{1}=\mu_{2}$ and $\mu_{12}=0$ for LRR implies that the positive solution of this equation has a simple expression,
@@
The pairs $(A,r_{1})$ determined by (118) for $\xi=1$ are shown in Figure 3. With $\xi=1$ and $r_{1}$ given by (119), the average sample sizes are bounded by (79) and (80).
Algorithm 1 (see Appendix A) describes the estimation procedure for LRR, as well as the estimator properties.
The simulation results for $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ with $\xi=1$ and varying $r_{1}$ are analogous to those shown for RR in Figure 4 (light dots), and are omitted for brevity. As in that case, with $r_{1}$ chosen as in (119) the bounds (79) and (80) are close to the actual average numbers of samples, and these approximately take their minimum values with respect to $r_{1}$. The simulated MSE, also omitted, has a similar behavior to that in Figure 6, with values always smaller than the target $A$, and very close to it unless $A$ is large.
Figure 12 shows, in two specific cases, the theoretical bounds (79) and (80) and simulated $\operatorname{E}[M_{i}+N_{i}]\phi$, for $\xi=1$ and with $r_{1}$ given by (119). The plotted values have a variation pattern similar to that observed for RR, with jump discontinuities and steep changes; again, the latter are only visible in the simulation results, and most apparent for large $A$. The values are slightly lower than those for RR in the rightmost region of the graphs (compare Figures 12(a) and 12(b) with 5(a) and 5(b) respectively), due to the fact that $\mu_{1}$ is $1$ for LRR and $2$ for RR. The effect of $\mu_{1}$ is only appreciable for large $A$, as it stems from (79) and (80). The ratio between average sample sizes obtained from simulation is close to $\lambda$, as can be seen from Figure 12 (note that the ratio between the bounds is exactly $\lambda$). More specifically, the achieved ratio deviates from $\lambda$ by small or very small percentages, very similar to those in RR (results not shown).
@@
As for the efficiency with element sampling, particularizing (87) for $\zeta=\Theta$, with $\partial\Theta/\partial p_{1}=1/p_{1}$, $\partial\Theta/\partial p_{2}=-1/p_{2}$ and substituting (79), (80) and (113) results in the same lower bound (89) as for RR (although the obtained values will be slightly different because $\mu_{1}$ and $r_{1}$ are), and the efficiency approaches $1$ when $A$ and $\phi$ tend to 0. Figure 13 compares this bound with simulation results, considering only two specific cases for brevity. The efficiency values from simulation are computed as explained in Section 2.1.4. Results are similar to those for RR (Figures 9(b) and 9(d)), except that for large $A$ the efficiency is larger in LRR, as is the difference between simulation and bound.
@@ 3.2 Group sampling
The estimation of LRR with group sampling uses, as in the RR case, a number of groups $G$ given by (91) in order to provide the necessary amounts of individual samples of the two populations. The procedure is the same as in Section 2.2.1: a new group is taken whenever a sample of either population is required and no surplus samples of that population are available from previous groups. At the end of the process, any leftover samples are discarded.
The analysis in Appendix B.1 shows that the approximation (94) is also valid for LRR; and $\operatorname{E}[G]$ is then given by (109), with $Y_{0}$ as in (98) (and with the values of $\mu_{1}$, $\mu_{2}$ corresponding to LRR). Simulation results for the average number of groups, omitted for brevity, are very similar to those for RR (Figure 10), except that for large $A$ they are slightly lower than in RR, for the same reason as with element sampling.
The efficiency with group sampling has the same approximate expression (112) as for RR, with $\operatorname{E}[G]$ computed as indicated above. The simulation results, shown in Figure 14, follow the same pattern observed for element sampling: the theoretical approximation is accurate when $A$ and $\phi$ are small, and conservative otherwise; and for large $A$ the efficiency deviates more from the theoretical curve than in RR (compare with Figure 11).
@@ 4.1 Element sampling
Several approaches are conceivable to estimate the OR $\psi$ defined in (2). One method that could be employed is to estimate $p_{1}/p_{2}$ and $(1-p_{2})/(1-p_{1})$ separately, treating each as a RR and using the two-stage procedure described in Section 2. With this approach, the MSE in the estimation of $\psi$ depends on the errors in the two RR estimations; and the problem is how to distribute the target MSE between these two components so as to approximately achieve a desired ratio of average sample sizes. To this end, another sampling stage could be introduced before the RR estimations, but that would complicate the process. A better approach, which only requires two stages and results in good estimation efficiency, is based on estimating the odds $p_{1}/(1-p_{1})$ and $(1-p_{2})/p_{2}$ separately, using in each case the method in Mendo [16, section 2]. This is detailed next.
@@
$\operatorname{E}[N_{1}\mid s_{1}]$ $=\frac{s_{1}}{p_{1}}+\frac{s_{1}-\alpha}{1-p_{1}}=\frac{s_{1}-\alpha p_{1}}{p_{1}(1-p_{1})}$ (120)
@@
Using the same ideas as for the RR estimator (Section 2.1.1), it can be seen that to approximately achieve a given ratio of average sample sizes, considering only the samples used by the second stage for the moment, $\operatorname{E}[s_{1}]/\operatorname{E}[s_{2}]$ should be roughly proportional to
@@
where $\epsilon_{1}=\epsilon_{2}=1/2$ as in Section 2. For a target relative MSE given by $A$, the second-stage IBS parameters $s_{1}$ and $s_{2}$ are obtained from $X$ by solving the equation system formed by (20) and (33), where the values of the design parameters $\gamma$, $\delta_{1}$ and $\delta_{2}$ are yet to be specified; and the solutions will then have to be rounded to integer values, as usual. Once $s_{1}$, $s_{2}$ are known, the second stage is carried out, from which $N^{\prime}_{1}$, $N^{\prime\prime}_{1}$, $N^{\prime}_{2}$, $N^{\prime\prime}_{2}$ are obtained. According to Mendo [16, section 2], and taking into account that $\alpha=2$,
@@
This is the case for OR estimation, and will also be true for LOR estimation (Section 5). Thus, in the sequel (141) will be assumed to hold. The two indicated conditions then reduce to
@@
An argument analogous to that used in Section 2.1.2 shows that (22) is satisfied.
The values of $s_{1}$, $s_{2}$ before rounding, obtained by solving (20) and (33) with $\delta_{1}$, $\delta_{2}$ and $\gamma$ as in (142)–(144), are given by (64)–(66) as usual. The curvature function is defined as in (76) with the terms $1/A+r_{1}+\mu_{1}+\xi$ replaced by $1/A+3r_{1}/2+\mu_{1}+\xi$ and with $\mu_{1}=\mu_{2}$:
@@
Similarly to previous estimators, $\xi$ is set to 11 and $r_{i}$ is chosen as in (81), where $c(A,r_{1},\xi)$ is given by (145). Equivalently, the curve formed by the pairs $(A,r_{1})$ that solve (146) can be plotted, as shown in Figure 3; and then, for a given $A$, (81) corresponds to rounding up the value obtained from the curve, with a minimum of $3$. Defining
@@
this choice of $\xi$ and $r_{1}$ results, by analogy with Section 2.1.3, in the upper bounds
@@
The estimation procedure for OR is specified in Algorithm 2 (see Appendix A), where the properties of the estimator are also indicated. (The algorithm covers the LOR case as well, to be presented in Section 5.)
The relative MSE obtained from simulations is compared with the target $A$ in Figure 15. The simulation consists, as with previous estimators, of $10^{6}$ realizations for each combination of parameters. The difference between simulation results and target is seen to increase with $A$ and with $\phi$, as for RR. However, in OR the difference observed for large $\phi$ vanishes when $A$ is small, unlike in RR (compare the leftmost parts of Figures 6(a) and 15(a)). This is related to the fact that the terms $1-\bar{p}_{1}/(s_{1}-2+2p_{1})$ and $1-\bar{p}_{2}/(s_{2}-2p_{2})$ in (129), which are replaced by $1$ in the uniform bound (130), approach $1$ for large $s_{1}$, $s_{2}$. Thus when $A$ is small, which gives large values of $s_{1}$ and $s_{2}$, the uniform bound (130) is almost as good as (129). In RR, on the other hand, the corresponding factors are $(s_{1}-2)(1-p_{1})/(s_{1}-2+2p_{1})<1-p_{1}$ and $1-p_{2}$, as is seen comparing (27) and (28), and these do not tend to $1$ for large $s_{1}$, $s_{2}$. (In LRR the factors are more cumbersome, see Mendo [16, theorem 2]; but for large $s_{1}$, $s_{2}$ they tend to the same values $1-p_{1}$, $1-p_{2}$ as in RR.)
@@
Figure 16 shows simulation results for $\operatorname{E}[M_{i}+N_{i}]\bar{\phi}$, $i=1,2$. The values are very similar to those for $\operatorname{E}[M_{i}+N_{i}]\phi$ in RR (Figures 5(a) and 5(b)) except that for OR they are somewhat larger in the rightmost part of the graphs. This is due to the $3/2$ factor in OR, whose effect on the average sample sizes is only significant for large $A$.
@@
Figure 17 shows simulation results for $\eta_{\mathrm{el}}$, and compares them with the bound (151). In the simulation, $\eta_{\mathrm{el}}$ is computed as in previous sections, i.e. using sample averages in (150). The values are seen to be similar to those for RR and LRR, except that with large $A$ the efficiency is slightly lower for OR. Again, this is a consequence of the $3/2$ factor. As in RR and LRR, the difference between simulation and bound tends to be larger when $\phi$ is increased. However, in OR that difference vanishes for small $A$, unlike in the other cases (compare Figure 17(a) with Figure 9(b) or 13(a)). This agrees with the behavior of the MSE discussed earlier.

### v3

@@ 2.1.3 Analysis of the approximation and bounds on average sample sizes
The ratio $\operatorname{E}[M_{1}+N_{1}]/\operatorname{E}[M_{2}+N_{2}]$ obtained from simulation is represented in Fig. 8. As can be seen, it is in general close to $\lambda$, and very close for small or moderate values of $A$. The relatively larger deviations in the rightmost region of the graphs can again be attributed to the fact that for large $A$, since $r_{1}$ is limited from below by $3$ in (81), the value of $r_{1}$ does not result in $c(A,r_{1},\xi)$ close to 0. In any case, the deviation between the actual ratio and $\lambda$ is less than $11\%$ for all combinations of parameters shown in the figure, and much lower for $A$ small.
@@ 2.1.4 Estimation efficiency
This matrix is readily computed from (83) as $F_{i,i}=n_{i}/(p_{i}(1-p_{i}))$, $F_{i,j}=0$ for $i\neq j$ (which reflects the fact that observations from one population give no information about the other), and therefore (84) becomes
@@
Figure 9 shows the efficiency $\eta_{\mathrm{el}}$ obtained from Monte Carlo simulation. Its value is computed using (88) with $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ replaced by the corresponding sample means and $\operatorname{Var}[\hat{\theta}$] replaced by the sample MSE. The bound (89) is also plotted. As seen in the figure, the efficiency is high for the values of the target $A$ commonly used in practice, and in particular it is close to $1$ for small $A$. For example, $A=0.04$, corresponding to a relative RMSE of $20\%$, gives values of $\eta_{\mathrm{el}}$ around $80\%$. The efficiency as a function of $A$ exhibits the same discontinuities and steep changes that have been identified for the relative MSE and for $\operatorname{E}[M_{i}+N_{i}]$. The simulation results deviate more from the theoretical bound for large $A$ (rightmost part of the graphs) or for large $\phi$ (Fig. 9(b)). This is explained by the fact that in these conditions the relative MSE is considerably smaller than the target, and the average sample sizes are also considerably smaller than their bounds, as discussed previously. On the other hand, the bound is quite tight for small or moderate values of $A$ and $\phi$, which is precisely when it is most important to characterize $\eta_{\mathrm{el}}$ accurately, as sample sizes are large in that case.
It is easily seen that $\lim_{A\rightarrow 0}Ar_{1}^{2}=1$ for $A$ and $r_{1}$ related by (82), and thus also for $r_{1}$ obtained from (81). Combining this result with (89), it follows that $\eta_{\mathrm{el}}$ tends to $1$ when $A$ and $\phi$ tend to 0. This is in consonance with the values shown in Fig. 9.
@@ 2.2.1 Estimation procedure
Estimation with group sampling thus proceeds as in Sect. 2.1, but with an added “outer layer” that translates group sampling into element sampling as described. With the above definition of $M_{i}$ and $N_{i}$, it follows that the number of required groups is
@@ 2.2.2 Average number of groups
and noting that $\max\{x,y\}=(x+y)/2+|x-y|/$2, it is possible to express $\operatorname{E}[G]$ as
@@
Using approximations (37) and (38) for $s_{1},s_{2}$ and including the rounding term $\xi$ as in Sect. 2.1.2 (this has very little effect on the approximations obtained here, but it is done for consistency), (95) becomes
@@
Figure 10 represents the normalized average number of groups obtained from simulation, as well as its theoretical approximation (109). Overall, the graphs follow similar patterns to those observed for $\operatorname{E}[M_{i}+N_{i}]\phi$ with element sampling. The approximation is accurate for small or moderate $A$, and less so for large $A$. Discontinuities due to the choice of $r_{1}$ are minimally visible, and steep changes due to rounding $s_{1}$ and $s_{2}$ can be observed in the simulation results for large $A$. Comparing $\operatorname{E}[G]$ for different values of $\theta$ only makes sense when $l_{1}$ and $l_{2}$ are fixed; that is, between Figs. 10(a) and 10(b) or between Figs. 10(c) and 10(d). For $l_{1}$ and $l_{2}$ given, it can be seen from (109) that the minimum of $\operatorname{E}[G]$ with respect to $\theta$ does not necessarily occur when $\lambda\theta=1$, as was the case for element sampling.
@@ 2.2.3 Estimation efficiency
The efficiency of the estimator is defined, as in Sect. 2.1.4, by comparing the estimation variance with the lowest variance that could be achieved by a fixed-size estimator with the same average numbers of samples, as given by the Cramér–Rao bound. With group sampling, the average number of samples required from population $i$ is $\operatorname{E}[G]l_{i}$, $i=1,2$. Equating $n_{i}$ to $\operatorname{E}[G]l_{i}$ in (86), the efficiency $\eta_{\mathrm{gr}}$ for an unbiased estimator of a generic parameter $\zeta$ with group sampling is obtained as
@@
Simulation results for the efficiency are shown in Fig. 11. The values are obtained using (111) with $\operatorname{E}[G]$ and $\operatorname{Var}[\hat{\theta}$] replaced by the sample mean and sample MSE. The approximation (112) is also displayed. The figure contains only two specific cases for brevity. The results are in general little sensitive to $\lambda$ and $\theta$; but, similarly to what was observed with element sampling, the theoretical approximation becomes more conservative for large $A$ or $\phi$. Comparing with Fig. 9, group sampling is seen to be less efficient than element sampling, in accordance with the number of required groups being the maximum over the two populations, as given by (91). This causes an efficiency loss of approximately $0.15$ for values of $A$ in the range 0.010.010.10.1, and a less substantial loss for $A$ small.
@@ 3 Estimation of log relative risk
This section describes the LRR estimator and analyzes its properties. The presentation, as will also be the case for subsequent estimators, follows the same logical course as in Sect. 2, but can be shorter, thanks to the similarities with RR. Section 3.1 addresses the estimation procedure with element sampling, and Sect. 3.2 considers group sampling.
@@ 3.1 Element sampling
$\operatorname{Var}[\hat{\Theta}^{2}]=\operatorname{E}[\operatorname{E}[\hat{\Theta}^{2}\mid s_{1},s_{2}]]-\Theta^{2}<\operatorname{E}\left[e(s_{1},s_{2})+\Theta^{2}\right]-\Theta^{2}\leq A.$ (117)
@@
As in the RR case, $s_{1}$ and $s_{2}$ are obtained from the equation system formed by (20) and (33), except with $\mu_{1},\mu_{2},\mu_{12}$ given by (116). The solution is (64)–(66). The values of $s_{1}$ and $s_{2}$ should then be rounded while satisfying (31). Using the first-order approximations (37)–(42), $\gamma$ and $\delta_{1}$ are obtained as in (61) and (62); and (22) holds. The average sample sizes are approximately given by (70) and (71).
The curvature function is defined in the same way as for RR, and the condition $c(A,r_{1},\xi)=0$ is expressed by (82). The fact that $\mu_{1}=\mu_{2}$ and $\mu_{12}=0$ for LRR implies that the positive solution of this equation has a simple expression,
@@
The pairs $(A,r_{1})$ determined by (118) for $\xi=1$ are shown in Fig. 3. With $\xi=1$ and $r_{1}$ given by (119), the average sample sizes are bounded by (79) and (80).
Algorithm 1 (see Appendix A) describes the estimation procedure for LRR, as well as the properties of the estimator.
The simulation results for $\operatorname{E}[M_{i}+N_{i}]\phi$, $i=1,2$ with $\xi=1$ and varying $r_{1}$ are analogous to those shown for RR in Fig. 4 (light dots), and are omitted for brevity. As in that case, with $r_{1}$ chosen as in (119) the bounds (79) and (80) are close to the actual average numbers of samples, and these approximately take their minimum values with respect to $r_{1}$. The simulated MSE, also omitted, has a similar behavior to that in Fig. 6, with values always smaller than the target $A$, and very close to it unless $A$ is large.
Figure 12 shows, in two specific cases, the theoretical bounds (79) and (80) and simulated $\operatorname{E}[M_{i}+N_{i}]\phi$ for $\xi=1$ and with $r_{1}$ given by (119). The plotted values have a variation pattern similar to that observed for RR (Sect. 2.1.3), with jump discontinuities and steep changes; again, the latter are only visible in the simulation results, and most apparent for large $A$. The values are slightly lower than those for RR in the rightmost region of the graphs (compare Figs. 12(a) and 12(b) with 5(a) and 5(b) respectively), due to the fact that $\mu_{1}$ is $1$ for LRR and $2$ for RR. The effect of $\mu_{1}$ is only appreciable for large $A$, as it stems from (79) and (80). The ratio between average sample sizes from simulation is close to $\lambda$, as can be seen in Fig. 12 (note that the ratio between the bounds is exactly $\lambda$). More specifically, the achieved ratio deviates from $\lambda$ by small or very small percentages, very similar to those in RR (results not shown).
@@
As for the efficiency with element sampling, particularizing (87) for $\zeta=\Theta$, with $\partial\Theta/\partial p_{1}=1/p_{1}$, $\partial\Theta/\partial p_{2}=-1/p_{2}$ and substituting (79), (80) and (113) results in the same lower bound (89) as for RR (although the obtained values will be slightly different because $\mu_{1}$ and $r_{1}$ are), and the efficiency approaches $1$ when $A$ and $\phi$ tend to 0. Figure 13 compares this bound with simulation results, considering only two specific cases for brevity. The efficiency values from simulation are computed as explained in Sect. 2.1.4. Results are similar to those for RR (Figs. 9(b) and 9(d)), except that for large $A$ the efficiency is larger in LRR, as is the difference between simulation and bound.
@@ 3.2 Group sampling
The estimation of LRR with group sampling uses, as in the RR case, a number of groups $G$ given by (91) in order to provide the necessary amounts of individual samples of the two populations. The procedure is the same as in Sect. 2.2.1: a new group is taken whenever a sample of either population is required and no surplus samples of that population are available from previous groups. At the end of the process, any leftover samples are discarded.
The analysis in Appendix B.1 shows that the approximation (94) is also valid for LRR; and $\operatorname{E}[G]$ is then given by (109), with $Y_{0}$ as in (98) (and with the values of $\mu_{1}$, $\mu_{2}$ corresponding to LRR). Simulation results for the average number of groups, omitted for brevity, are very similar to those for RR (Fig. 10), except that for large $A$ they are slightly lower than in RR, for the same reason as with element sampling.
The efficiency with group sampling has the same approximate expression (112) as for RR, with $\operatorname{E}[G]$ computed as indicated above. The simulation results, shown in Fig. 14, follow the same pattern observed for element sampling: the theoretical approximation is accurate when $A$ and $\phi$ are small, and conservative otherwise; and for large $A$ the efficiency deviates more from the theoretical curve than in RR (compare with Fig. 11).
@@ 4.1 Element sampling
Several approaches are conceivable to estimate the OR $\psi$ defined in (2). One method that could be employed is to estimate $p_{1}/p_{2}$ and $(1-p_{2})/(1-p_{1})$ separately, treating each as a RR and using the two-stage procedure described in Sect. 2. With this approach, the MSE in the estimation of $\psi$ depends on the errors in the two RR estimations; and the problem is how to distribute the target MSE between these two components so as to approximately achieve a desired ratio of average sample sizes. To this end, another sampling stage could be introduced before the RR estimations, but that would complicate the process. A better approach, which only requires two stages and results in good estimation efficiency, is based on estimating the odds $p_{1}/(1-p_{1})$ and $(1-p_{2})/p_{2}$ separately, using in each case the method in Mendo [16, section 2]. This is detailed next.
@@
$\operatorname{E}[N_{1}\mid s_{1}]$ $=\frac{s_{1}}{p_{1}}+\frac{s_{1}-\alpha}{1-p_{1}}=\frac{s_{1}-\alpha p_{1}}{p_{1}(1-p_{1})},$ (120)
@@
Using the same ideas as for the RR estimator (Sect. 2.1.1), it can be seen that to approximately achieve a given ratio of average sample sizes, considering only the samples used by the second stage for the moment, $\operatorname{E}[s_{1}]/\operatorname{E}[s_{2}]$ should be roughly proportional to
@@
where $\epsilon_{1}=\epsilon_{2}=1/2$ as in Sect. 2. For a target relative MSE given by $A$, the second-stage IBS parameters $s_{1}$ and $s_{2}$ are obtained from $X$ by solving the equation system formed by (20) and (33), where the values of the design parameters $\gamma$, $\delta_{1}$ and $\delta_{2}$ are yet to be specified; and the solutions will then have to be rounded to integer values, as usual. Once $s_{1}$, $s_{2}$ are known, the second stage is carried out, from which $N^{\prime}_{1}$, $N^{\prime\prime}_{1}$, $N^{\prime}_{2}$, $N^{\prime\prime}_{2}$ are obtained. According to Mendo [16, section 2], and taking into account that $\alpha=2$,
@@
This is the case for OR estimation, and will also be true for LOR estimation (Sect. 5). Thus, in the sequel (141) will be assumed to hold. The two indicated conditions then reduce to
@@
An argument analogous to that used in Sect. 2.1.2 shows that (22) is satisfied.
The values of $s_{1}$, $s_{2}$ before rounding, obtained by solving (20) and (33) with $\delta_{1}$, $\delta_{2}$ and $\gamma$ as in (142)–(144), are again given by (64)–(66). The curvature function is defined as in (76) with the terms $1/A+r_{1}+\mu_{1}+\xi$ replaced by $1/A+3r_{1}/2+\mu_{1}+\xi$ and with $\mu_{1}=\mu_{2}$:
@@
Similarly to previous estimators, $\xi$ is set to 11 and $r_{i}$ is chosen as in (81), where $c(A,r_{1},\xi)$ is given by (145). Equivalently, the curve formed by the pairs $(A,r_{1})$ that solve (146) can be plotted, as shown in Fig. 3; and then, for a given $A$, (81) corresponds to rounding up the value obtained from the curve, with a minimum of $3$. Defining
@@
this choice of $\xi$ and $r_{1}$ results, by analogy with Sect. 2.1.3, in the upper bounds
@@
The estimation procedure for OR is specified in Algorithm 2 (see Appendix A), where the properties of the estimator are also indicated. (The algorithm covers the LOR case as well, to be presented in Sect. 5.)
The relative MSE obtained from simulations is compared with the target $A$ in Fig. 15. The simulation consists, as with previous estimators, of $10^{6}$ realizations for each combination of parameters. The difference between simulation results and target is seen to increase with $A$ and with $\phi$, as for RR. However, in OR the difference observed for large $\phi$ vanishes when $A$ is small, unlike in RR (compare the leftmost parts of Figs. 6(a) and 15(a)). This is related to the fact that the factors $1-\bar{p}_{1}/(s_{1}-2+2p_{1})$ and $1-\bar{p}_{2}/(s_{2}-2p_{2})$ in (129), which are replaced by $1$ in the uniform bound (130), approach $1$ for large $s_{1}$, $s_{2}$. Thus when $A$ is small, which gives large values of $s_{1}$ and $s_{2}$, the uniform bound (130) is almost as good as (129). In RR, on the other hand, the corresponding factors are $(s_{1}-2)(1-p_{1})/(s_{1}-2+2p_{1})<1-p_{1}$ and $1-p_{2}$, as is seen comparing (27) and (28), and these do not tend to $1$ for large $s_{1}$, $s_{2}$. (In LRR the factors are more cumbersome, see Mendo [16, theorem 2]; but for large $s_{1}$, $s_{2}$ they tend to the same values $1-p_{1}$, $1-p_{2}$ as in RR.)
@@
Figure 16 shows simulation results for $\operatorname{E}[M_{i}+N_{i}]\bar{\phi}$, $i=1,2$. The values are very similar to those for $\operatorname{E}[M_{i}+N_{i}]\phi$ in RR (Figs. 5(a) and 5(b)) except that for OR they are somewhat larger in the rightmost part of the graphs. This is due to the $3/2$ factor in OR, whose effect on the average sample sizes is only significant for large $A$.
@@
Figure 17 shows simulation results for $\eta_{\mathrm{el}}$, and compares them with the bound (151). In the simulation, $\eta_{\mathrm{el}}$ is computed as in previous sections, i.e. using sample averages in (150). The values are seen to be similar to those for RR and LRR, except that with large $A$ the efficiency is slightly lower for OR. Again, this is a consequence of the $3/2$ factor. As in RR and LRR, the difference between simulation and bound tends to be larger when $\phi$ is increased. However, in OR that difference vanishes for small $A$, unlike in the other cases (compare Fig. 17(a) with Fig. 9(b) or 13(a)). This agrees with the behavior of the MSE discussed earlier.